40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jun 25 09:58:08 2013, Erica, Notes, fiber optic, mode matching refcav_mm.JPG
    Reply  Mon Jul 8 21:52:04 2013, tara, DailyProgress, optic, mode matching to refcavs 
       Reply  Tue Jul 9 19:35:20 2013, tara, DailyProgress, optic, beat board is back 
          Reply  Thu Jul 11 00:19:52 2013, tara, DailyProgress, optic, both cavities are locked photo_2013_07_10.JPG
Message ID: 1231     Entry time: Tue Jul 9 19:35:20 2013     In reply to: 1228     Reply to this: 1233
Author: tara 
Type: DailyProgress 
Category: optic 
Subject: beat board is back 

I installed the beat board back behind the cavities. I still have not finished aligning both beams to the 1811.

  • Note about ACAV ( this path has PMC on it): After new mode matching with more visibility (from 80% to 95%), I can increase more gain and the error noise is getting lower. However, there is a problem with the beam reflected from the window of the tank. It overlaps with the main beam and cannot be blocked. I think this is the reason why we cannot suppress the error noise down to what we had before. I still need to convert the error noise back to frequency noise to see if it is below the estimated coating noise or not. If not, we have to reopen the chamber and tilt the cavity a bit. Rcav does not have this problem, the back reflection is away from the main beam and can be dump properly.
  • Note about RCAV: Erica and I plan to finish the EOM driver test tomorrow. After that I'll use it to drive the broadband EOM for locking RCAV to the cavity. The plan is to use one marconi to drive two EOM at the same frequency (14.75 MHz). We use a 4-way splitter for 2 EOM and 2 demodulations. I don't know how using same frequency for EOM will turn out (cross talk problem), but I want to see the first beat measurement within this week.
  • Note about beat setup: Evan calculated the mode matcing for beat setup, but I had to modify it. The first lenses were moved out of the board and mounted between the vacuum tank and the board due to space limitation. This might add some extra resonant peaks in the beat setup due to the long posts for lenses. The spot diameter on the PD is about 130um, which should be fine because 1811's diameter is ~300 um.
ELOG V3.1.3-